# EffLiFe: Efficient Light Field Generation via Hierarchical Sparse Gradient Descent

### v1

@@ 3.1 Preliminaries
3.2 Data Preparation for Optimization
As depicted in Fig. 3, our primary optimization process, Sparse Gradient Descent, relies on two inputs: filtered plane sweep volumes (PSVs) and an MPI. In the following section, we provide a brief overview of how these inputs are constructed and generated. Additionally, we discuss the occlusion filter module in detail in Sec. 3.4.
Hierarchical PSV Construction. Given $N$ source view images ${I}_{i=1}^{N}$ of size $H\times W\times 3$ and their corresponding camera poses $[K_{i},R_{i}|t_{i}]$, as well as for the reference view $I_{r}$, we construct the plane sweep volume $P$ of size $N\times D\times H\times W\times 3$ leveraging homographic warping:
$\mathcal{H}_{i}(d)=K_{i}R_{i}(\mathcal{I}+\frac{(R_{i}^{-1}t_{i}-R_{r}^{-1}t_{r})n^{T}R_{r}}{d})R_{r}^{-1}K_{r}^{-1},$ (6)
where $n$ denotes the plane normal and $\mathcal{I}$ is an identity matrix. And the homography matrix $\mathcal{H}_{i}(d)$ warps a pixel $(u,v)$ of the reference view at depth $d$ to the source view $I_{i}$. The warped source view $I_{i}^{w}$ is defined as:
$I_{i}^{w}(u,v,d)=I_{i}(\mathcal{H}_{i}(d)[u,v,1]^{T}).$ (7)
By concatenating all the warped source view images $I_{i}^{w}$ along the depth dimension, the plane sweep volume $P$ is constructed. Before feeding $P$ into the network, we downsample it hierarchically at multiple resolutions. Specifically, we create the plane sweep volume $P_{m}$ of size $N\times D\times H/2^{m}\times W/2^{m}\times 3$ for $m$’s iteration. In the default implementation, the number of iterations is set to 3.
Initial MPI Generation. A 3D neural network accepts the reshaped coarsest Plane Sweep Volume (PSV) of dimension $3N\times D\times H\times W$ as input and produces the Multi-plane Image (MPI) at the lowest resolution. The neural network is trained to aggregate information across images and generate a coarse MPI for subsequent refinement.
@@ 3.3 Hierarchical Sparse Gradient Descent
Hierarchical sparse gradient descent is a coarse-to-fine process that progressively refines the Multi-plane Image (MPI) using MPI gradients. Among these gradients, the one relative to the color channel holds utmost significance as it provides the network with essential visibility information, commonly referred to as the ”alpha gradient”. Fig. 4 (b) visually depicts that the alpha gradients for the majority of pixels are close to zero. These pixels have minimal influence on the final rendering results; however, they consume significant computational resources during the optimization process. This observation shows significant potential for sparse optimization and acceleration.
@@
So we propose sparse gradient descent, an adaptive approach that concentrates on important MPI gradients for efficient optimization. Specifically, this method consists of three steps: gradient formulation, sparse gradient selection, and sparse gradient update. The detailed process is shown in Fig. 5.
@@
Gradient Formulation. To obtain the sparse MPI gradient, we first formulate the input gradient components for learned gradient descent. They are defined as the concatenation of the input PSV $P$, the input MPI $M$ and the warped alpha gradient $\mathcal{A}^{w}_{i}$ of each source view. To compute the warped alpha gradient for a single source view, the MPI of the reference view is initially warped to this source view using homographic warping (Eq. 6) to obtain $M_{i}$. Then the alpha gradient $\mathcal{A}_{i_{d}}$ corresponding to depth $d$ is calculated by the following equation:
$\mathcal{A}_{i_{d}}=\frac{\partial\mathcal{O}(M_{i})}{\partial c_{d}}=\alpha_{d}\prod_{j=d+1}^{D}(1-\alpha_{j}).$ (8)
@@
After this, the warped alpha gradient $\mathcal{A}^{w}_{i}$ is derived by applying inverse homographic warping Eq. 6 from the source view to the reference view. So the final gradient components $V$ are defined as:
$V=concat(P,M,\mathcal{A}^{w}_{1},...,\mathcal{A}^{w}_{N}),$ (9)
where $P$ serves as the raw input to provide details of higher resolution, the MPI $M$ is the basis for optimization, and alpha gradient $\mathcal{A}^{w}_{i}$ provides useful visibility information to eliminate occlusion.
Sparse Gradient Selection. As shown in Fig. 5 (a), $V$ is a large volume that would require significant computational resources. Therefore, we employ sparsification on $V$ utilizing the alpha gradient $\mathcal{A}$ of the MPI, resulting in a sparsified version denoted as $V_{s}$. Specifically, the selection process involves choosing only the $k$ pixels with the highest alpha gradient along the $D$ axis, resulting in the compression of the original input volume $V$ from $C\times D\times H\times W$ to $C\times k\times H\times W$. To strike a balance between rendering quality and efficiency, $k$ is set for 5 for the default model. It is important to note that we choose alpha gradient $\mathcal{A}$ rather than $\alpha$ as the sparsification criterion for two reasons: 1) The alpha values of the farthest few planes in a generated MPI are close to 1, as they represent relatively static background content. If pixels were selected based on $\alpha$, the majority would be from the background rather than the object surfaces; 2) Alpha gradient $\mathcal{A}$ can be regarded as the weight of each color plane $c_{d}$ to the final color, suggesting that pixels with higher alpha gradient contribute more to the final color.
The sparse indices $S$, which store the positions of the $k$ selected pixels in the original $D$ pixels, are also retained to facilitate the restoration of the compressed MPI residual. This allows it to be added to the input MPI for an update.
@@
Sparse Gradient Update. As shown in Fig. 5 (c), the sparsified input volume $V_{s}$ is fed to a 3D CNN, functioning as the learned gradient descent network, to derive the sparse MPI gradient for an MPI refinement iteration. The pixels in the sparse MPI gradient are then restored to their original positions using the sparse indices $S$, and subsequently added to the input MPI to obtain a refined version:
@@
where $\mathcal{N}_{w}$ is a 3D CNN that functions as the learend gradient descent neural network, and $\mathcal{R}$ is the operator that returned the pixels of the sparse MPI gradient to their original positions.
3.4 Occlusion-aware Iterative Refinement
Figure 6: The occlusion-aware iterative refinement module. Depth maps of the reference view and source views are rendered from the input MPI, which are then used to formulate the depth PSV using homowarping. For a pixel located in a specific depth plane of the depth PSV, the occlusion view index (source view 4) is determined based on a significant discrepancy between its rendered depth value and that of the reference view. Subsequently, the occlusion content is substituted with the mean value of the non-occlusion views (the source views 1-3).
The hierarchical sparse gradient descent strategy achieves significant acceleration. Nonetheless, ghosting effects at object boundaries may be observed in certain scenes with severe occlusion. The occlusion-aware iterative refinement module, illustrated in Fig. 6, seeks to eliminate occlusion within the input plane sweep volume $P$, employing the MPI $M$ from the previous iteration.
Specifically, the depth maps of the reference view and source views are rendered in the same manner as RGB images:
$\mathcal{O}_{d}(M)=\sum_{i=1}^{D}d_{i}\alpha_{i}\prod_{j=i+1}^{D}(1-\alpha_{j}),$ (11)
where $\mathcal{O}_{d}$ is the over composite operator on depth, and $d_{i}$ is the depth value of plane $i$. Utilizing homographic warping in Eq. 6, the depth plane sweep volume (DPSV) $P_{D}\in N\times D\times H\times W$ is generated by just replacing the RGB plane with the rendered depth planes. $P_{D}$ can be perceived as a 3D grid in the shape of $D\times H\times W$, with each voxel of the grid containing $N$ depth values corresponding to $N$ source views. By comparing $N$ depths at coordinate $(d,h,w)$ to the depth of the target view at coordinate $(h,w)$ where $d\in[0,D),h\in[0,H),w\in[0,W)$, the index of the occlusion view at $(d,h,w)$ can be readily identified if:
$\frac{1}{\mathcal{D}_{t}[h,w]}-\frac{1}{P_{D}[i,d,h,w]}>\epsilon,$ (12)
where $\mathcal{D}_{t}\in H\times W$ is the rendered depth map of the target view $t$ and $\epsilon$ is a threshold controlling the minimum tolerance for considering view $i$ at coordinate $(d,h,w)$ as the occlusion view. Next, the RGB color of source view $i$ at $(d,h,w)$ is substituted with the mean value calculated from the remaining source views. This adjustment enhances the visual coherence of all source views at $(d,h,w)$. By replacing the colors, the method achieves sharper edge rendering by eliminating sudden color discrepancies at object boundaries. Additionally, this operation does not disrupt areas of smooth depth transitions because small depth difference between source views and the target view is filtered by $\epsilon$.
3.5 Implementation Details
Network Details. The trainable components of the entire pipeline are the 3D CNNs that convert the input plane sweep volume into a multi-plane image in Fig. 3, and the one that refines the sparse MPI gradient in Fig. 5. In consideration of time efficiency, we use only a 3-layer 3D convolution block denoted in both figures. The kernel size and padding size of the first two convolution layers are 3 and 1, respectively, followed by a batchnorm layer and a relu layer. The third layer is a straightforward 1x1x1 3D convolution. For the ”doubled conv layers” setting of the ablation study, the 3D CNN is replaced with a stronger one that has doubled convolution layers and more channels, resulting in even higher visual metrics.
The backbone 3D CNN of the default model EffLiFe is shown in Fig. 7(a), and the backbone 3D CNN of EffLiFe-dc that has doubled convolution layers is shown in Fig. 7(b).
Figure 7: (a) The backbone 3D CNN of the default model EffLiFe, which contains only three convolution layers. (b) The backbone 3D CNN of EffLiFe-dc, which contains six convolution layers.
@@ 3.5 Implementation Details
Training Loss. We train our network using synthesized views as supervision and choose the commonly used deep feature matching loss $L_{VGG}$[37] as the basic loss function. In detail, our overall rendering loss $L_{r}$ is a weighted average loss of all iterations, and the rendering loss for each iteration is also a weighted average one for both the synthesized reference and source views:
$L_{r}=\sum_{i=1}^{m}\lambda_{i}(L_{VGG}(I_{r_{i}},\tilde{I}_{r_{i}})+\mu\sum_{j=1}^{N}L_{VGG}(I_{j_{i}},\tilde{I_{j_{i}}})),$ (13)
@@
where $I_{r_{i}}$ and $I_{j_{i}}$ are the ground truth reference and source images at iteration $i$, $\tilde{I}_{r_{i}}$ and $\tilde{I}_{j_{i}}$ are the rendered reference and source images from the generated MPI at iteration $i$. And $m$ is the number of iterations, $N$ is the number of input source images, $\lambda_{i}$ is a hyper-parameter weighing the importance of each iteration, and $\mu$ is also a hyper-parameter balancing reference and source image supervision.
@@
In addition, we propose a sparsity loss $L_{s}$ that regularizes the alpha values of the MPI pixels to be close to 0 or 1:
@@
where $\alpha_{h,w,d}$ is the alpha channel of the MPI, and $H,W,D$ are the height, width and number of depth planes of the volume. This loss aims to reduce the entropy of the MPI with respect to its $\alpha$ channel. It is ideal for sparse gradient selection, as it allows us to select more informative top $k$ samples that lie around the surfaces of a scene. The overall training loss is:

### v3

@@ 3.1 Preliminaries
3.2 Initial MPI Generation
In this section, we delve into generation of the Initial MPI, a pivotal component in our optimization process, Hierarchical Sparse Gradient Descent. This process involves both the construction of the PSV and a straightforward network forwarding.
Hierarchical PSV Construction. Given $N$ source view images ${I}_{i=1}^{N}$ of size $H\times W\times 3$ and their corresponding camera poses $[K_{i},R_{i}|t_{i}]$, as well as for the reference view $I_{r}$, we construct the plane sweep volume $P$ leveraging inverse homographic warping:
$\mathcal{H}^{-1}_{i}(d)=K_{i}R_{i}(\mathcal{I}+\frac{(R_{i}^{-1}t_{i}-R_{r}^{-1}t_{r})n^{T}R_{r}}{d})R_{r}^{-1}K_{r}^{-1},$ (6)
where $n$ denotes the plane normal and $\mathcal{I}$ is an identity matrix. The inverse homography matrix $\mathcal{H}^{-1}_{i}(d)$ warps the reference view meshgrid, defined by pixel positions $(u,v)$ at depth $d$, to each source view $I_{i}$. And the inverse warped source view $I_{i}^{w^{-1}}$ is grid-sampled by the inverse warped meshgrid:
$I_{i}^{w^{-1}}(u,v,d)=grid\_sample(I_{i},\mathcal{H}^{-1}_{i}(d)[u,v,1]^{T}),$ (7)
By concatenating all the inverse warped source view images $I_{i}^{w^{-1}}$ along the depth dimension, the plane sweep volume $P$ of size $N\times D\times H\times W\times 3$ is constructed.
$P=[I_{0}^{w^{-1}},I_{1}^{w^{-1}},...,I_{N}^{w^{-1}}]$ (8)
To enable optimization at different scales, we downsample it hierarchically at multiple resolutions. Specifically, we create the plane sweep volumes $\{P_{l}|1<=l<=L\}$ of size $N\times D\times H/2^{l}\times W/2^{l}\times 3$ for $L$ iterations.
Network Forwarding. The coarsest-level PSV is reshaped into $3N\times D\times H/2^{l}\times W/2^{l}$ and fed to a 3D neural network to produce the initial MPI. The neural network is trained to aggregate information across views and generate a coarse MPI for subsequent optimization.
@@ 3.3 Hierarchical Sparse Gradient Descent
Hierarchical sparse gradient descent is a coarse-to-fine approach that progressively refines the MPI using sparse gradients. This design is supported by two key insights: Coarse-to-fine Appearance Refinement and Sparse Gradients Suffice for Light Field Reconstruction. The coarse-to-fine geometry refinement has previously demonstrated its efficacy in various MVS works[24, 25, 26, 27, 28, 29]. In this context, a low-resolution geometric scaffold can be rapidly derived and employed as a prior to guide the subsequent high-resolution detail refinement process. Similarly, the task of deriving an MPI involves allocating scene contents to their respective depth planes. So the hierarchical pipeline is well-suited for MPI generation, offering a structured foundation upon which the learned gradient descent algorithm can iteratively refine appearance details. And the choice of sparse gradients stems from the empirical observation in Fig. 6 that after only one iteration of network forwarding, above half of the voxels in an MPI become empty. Hence, any subsequent optimization targeting these empty regions does not lead to appreciable enhancements in rendering quality but instead incurs unnecessary computational overhead.
Figure 6: Proportions of empty pixels of individual MPI planes throughout optimization iterations. (a) A rendered image of RealLife (without using HSGD) with 40 MPI planes. (b) A subset of MPI planes (i.e., the 12th, 16th, and 20th MPI planes) of three iterations. (c) Proportions of empty pixels in individual MPI planes. (pixels characterized by $\alpha<0.1$ are considered empty, as their contribution to rendering quality is marginal.)
@@
So we propose Hierarchical Sparse Gradient Descent, an adaptive and iterative approach that only leverages sparse gradients for efficient optimization. Specifically, this method consists of three steps: gradient formulation, gradient sparsification, and sparse gradient update. The detailed process is shown in Fig. 5.
Gradient Formulation. To obtain the sparse MPI gradient, we first formulate the components of input gradients. They are defined as the concatenation of the input PSV $P$ and the inverse warped alpha gradient $\mathcal{A}^{w^{-1}}_{i}$ of each source view $i$. Theoretically, the full components of MPI gradients[1] should be in the format of Eq. 9 (The detailed derivation can be referred from [1].):
$\nabla=[P,\widetilde{P},\mathcal{A}^{w^{-1}}_{1},...,\mathcal{A}^{w^{-1}}_{N},\mathcal{T}^{w^{-1}}_{1},...,\mathcal{T}^{w^{-1}}_{N}],$ (9)
where $\widetilde{P}$ is the PSV constructed with rendered images at each source view from the generated MPI, and $\mathcal{T}^{w^{-1}}_{i}$ is the warped MPI gradient respective to MPI’s $\alpha$ channel. But we simplify the gradients to be in the format of 10:
~$\widetilde{\nabla}=[P,\mathcal{A}^{w^{-1}}_{1},...,\mathcal{A}^{w^{-1}}_{N}],$ (10)
for several reasons:
1) The iterative optimization resembles a residual refinement network, where the PSV $P$ is intuitively effective stable and precise residual information directly derived from the input data.
2) The MPI generation process involves the allocation of scene contents to their respective depth planes, with $\mathcal{A}$ being in perfect alignment with the scene contents. This alignment constitutes a crucial gradient component for the accurate construction of MPIs.
3) In the quest for an optimal trade-off between rendering quality and computational efficiency, it is proved in DeepView[1] that $\widetilde{P}$ and $\mathcal{T}$ can only marginally enhance the ultimate rendering quality, but their large spatial sizes can potentially hinder real-time efficiency.
@@
To compute $\mathcal{A}^{w_{i}^{-1}}$ for a single source view ${i}$, the MPI of the reference view is initially warped to source view $i$ to obtain $M_{i}$. Then the alpha gradient $\mathcal{A}_{i_{d}}$ corresponding to depth $d$ is calculated by the following equation:
$\mathcal{A}_{i_{d}}=\frac{\partial\mathcal{O}(M_{i})}{\partial c_{d}}=\alpha_{d}\prod_{j=d+1}^{D}(1-\alpha_{j}),$ (11)
@@
After this, the inverse warped alpha gradient $\mathcal{A}^{w_{i}^{-1}}$ is derived by applying inverse homographic warping Eq. 6 from the source view to the reference view.
After formulating the gradients, the full 3D volume $V$ to be sparsified and optimized is the concatenation of $\widetilde{\nabla}$ and $M$:
$V=[\widetilde{\nabla},M],$ (12)
Gradient Sparsification. To safely sparsify gradient components without adversely affecting the final rendering quality, it becomes imperative to establish criteria for the exclusion of voxels that impart minimal contributions. A commonly adopted strategy in various Multiview Stereo (MVS) studies[24, 25, 26] is to reduce the number of depth layers in the full cost volume from $D$ to $k$, where $k$ is usually an odd number, composed of the plane with the highest probability and its adjacent $22\times(k-1)/$ 2 layers. However, this MVS-sampling strategy primarily relies on surface-centric geometric considerations, which might be too aggressive to capture some light field details like semi-transparent objects and thin plates. If we revisit Eq. 2, it can be rewritten as:
$\mathcal{O}(M)=\sum_{d=1}^{D}c_{d}\mathcal{A}_{d},$ (13)
from which we can easily find that the influence of the color component $c_{d}$ of each MPI plane on the ultimate color is determined by the alpha gradient derived from the reference view $\mathcal{A}_{d}$. The alpha gradient aligns more closely with the scene contents, as it takes into consideration not only solid geometry but also semi-transparent surfaces and thin plates, making it a more comprehensive and effective choice. In the ablation study of Sec. 5.5, we demonstrate that our sparsification strategy consistently outperforms the MVS-sampling approach in terms of overall performance. Consequently, the criterion for gradient sparsification is determined by $\mathcal{A}_{d}$.
The remaining challenge is to decide the optimal $k$ for gradient sparsification. Fig. 7 provides insights into the determination of the optimal value for $k$. For instance, when $k=7$, the rendered color almost recovers more than $80\%$ of the original color, indicating a trade-off between computational efficiency and rendering quality. Generally, it appears from Fig. 7 that values of $k$ in the range of 5 to 7 represent favorable choices, achieving a balance between time efficiency and rendering quality. Further ablation experiments on different values of $k$ will be presented in Sec. 5.5.
Figure 7: The influence of $k$ on the rendering quality. (a) A rendered image of RealLiFe (without using HSGD) with 40 MPI planes. (b) Top 5 MPI planes with the highest alpha gradients $\mathcal{A}$ for the red pixel. (The MPI is multiplied by $\mathcal{A}$ to better visualize its contribution to the final rendering result.) (c) The alpha gradients of the red pixel across 40 MPI planes, with the red dotted lines partitioning $d$ based on whether its alpha gradient falls within the top $k$. (d) The color ratio ($RGB_{k=3,5,7}/RGB_{k=40}$) and PSNR of the rendered red pixel in comparison to when $k=40$.
After sparsification, only $k$ voxels with the highest alpha gradient along the depth axis are kept, resulting in the compression of the input volume $V$, which is the concatenation of the original gradient volume $\widetilde{\nabla}$ and the input MPI $M$, from $C\times D\times H\times W$ to $C\times k\times H\times W$. The sparse indices $S$, which store the positions of the $k$ selected voxels in the original depth axis, are also retained to enable the generated spasified MPI residual to be added to the input complete MPI for an update. It is evident that while the complete spatial structure of the light field is sacrificed after sparsification, the sparsified layers are retained in their original spatial order, preserving the potential surface structure of the light field for subsequent optimization.
Sparse Gradient Update. After the gradient components of each voxel has been decided, a network backbone needs to be decided for the learned gradient descent module. A CNN with large perceptive fields is proved effective in DeepView[1]. However, after gradient sparsification, an issue arises where adjacent voxels may not necessarily belong to the same depth plane. This raises concerns about the suitability of convolutional operations on these voxels. So a straightforward decision for the backbone would be an MLP, which treats each voxel independently. However, this choice compromises the benefits of an extensive receptive field, resulting in suboptimal performance. Moreover, even though the depths of a selected subset of voxels may exhibit considerable disparities, convolutional processing enables the model to adapt to these variations by learning shared features and patterns across different regions. We further prove the adaptability of the 3D CNN in enhancing rendering quality without introducing conspicuous artifacts through qualitative analysis in Sec. 10 and quantitative evaluation in Sec. 5.2.
@@
As shown in Fig. 5 (c), the sparsified input volume $V_{s}$ goes through a 3D CNN, functioning as the learned gradient descent network, to derive the sparse MPI residual for an MPI refinement iteration. The voxels in the sparse MPI residual are then restored to their original positions using the sparse indices $S$, and subsequently added to the input MPI to obtain a refined version:
@@
where $\mathcal{N}_{w}$ is the 3D CNN that functions as the learend gradient descent neural network, and $\mathcal{R}$ is the operator that restored the voxels of the sparse MPI residual to their original positions.
3.4 Training
@@
During training, the network parameters are progressively optimized by minimizing the difference between the synthesized views and the ground truth novel views. And we choose the commonly used deep feature matching loss $L_{VGG}$[46] as the basic loss function. In detail, our overall rendering loss $L_{r}$ is a weighted average loss of all iterations, and the rendering loss for each iteration is also a weighted average one for both the synthesized reference and source views:
$L_{r}=\sum_{i=1}^{l}\lambda_{i}(L_{VGG}(I_{r_{i}},\tilde{I}_{r_{i}})+\mu\sum_{j=1}^{N}L_{VGG}(I_{j_{i}},\tilde{I_{j_{i}}})),$ (15)
@@
where $I_{r_{i}}$ and $I_{j_{i}}$ are the ground truth reference and source images at iteration $i$, $\tilde{I}_{r_{i}}$ and $\tilde{I}_{j_{i}}$ are the rendered reference and source images from the generated MPI at iteration $i$. And $l$ is the number of iterations, $N$ is the number of input source images, $\lambda_{i}$ is a hyper-parameter weighing the importance of each iteration, and $\mu$ is also a hyper-parameter balancing reference and source image supervision.
@@
In addition, we propose a sparsity loss $L_{s}$ that regularizes the alpha values of the MPI to be close to 0 or 1:
@@
where $\alpha_{h,w,d}$ is the alpha channel of the MPI, and $H,W,D$ are the height, width and number of depth planes of the volume. This loss aims to reduce the entropy of the MPI with respect to its $\alpha$ channel. It is ideal for gradient sparsification, as it allows us to select more informative top $k$ samples that lie around the surfaces of a scene. The overall training loss is:
@@
4 Implementation Details
The generalized model was trained using an RTX 3090 GPU with the Adam optimizer[47]. The initial learning rate was configured at $31e-$ 3, with a reduction by half if the training loss consistently decreased for a continuous span of 10 epochs. The model underwent training for 1000k iterations, taking approximately 13 hours to complete. The trainable components within the entire pipeline included the 3D CNNs responsible for converting the input plane sweep volume into a multi-plane image (as depicted in Fig. 4) and refining the sparse MPI gradient (as illustrated in Fig. 5). The structure of the 3D CNN, as shown in Fig. 8, comprised 5 basic convolutional blocks (a convolution layer followed by a batchnorm layer and a relu layer) and 1 1x1x1 3D convolution layer. To further improve the efficiency of our method, we optimize to run our pipeline with TensorRT [^1] https://developer.nvidia.com/tensorrt.
To configure the loss function hyperparameters, a preference was given to increased color supervision from the reference view and the last iteration, while maintaining the impact of the sparsity loss. An example set of these parameters was as follows: $\lambda_{1}=0.2$, $\lambda_{2}=0.3$, $\lambda_{3}=0.5$, $\mu=0.5$, $\lambda_{s}=0.2$. Regarding experimental parameters, the model utilized 40 MPI planes, underwent 3 network iterations, and employed a sparsification factor $k$ set to 5 as the default. These settings may vary in the context of ablation experiments.
Figure 8: The backbone 3D CNN of RealLiFe, which contains only three convolution layers.